OpenStudy (anonymous):

How many moles of HNO3 will be produced when .65 grams of N2O5 reacts?

OpenStudy (anonymous):

You need a balanced chemical equation for this. Does it specify what the N205 reacts with?

OpenStudy (anonymous):

N2O5 = H2O -> 2HNO3

OpenStudy (anonymous):

Okay, this is actually really simple. All you need is the mole/mole ratio (which you have in the equation) and the molar mass of each substance, which you get from a periodic table. Find the mass of N2O5, which is 108.01 g/mol, and HNO3, which is 63.01. Set up the equation \[0.65g N _{2}O _{5}^{?}(\frac{ 1 mol N _{2} O _{5}}{ 108.01 g })(\frac{ 2 mol HNO _{3} }{ 1 mol N _{2} O _{5} })(\frac{ 63.01 g HNO _{3} }{ 1 mol HNO _{3}}) = \] and solve.

OpenStudy (anonymous):

Oops, ! accidentally added grams of HNO3. Forget that last parenthesis section and you should get 0.012035922 mol HNO3, which you would probably only keep two significant figures of, so 0.012 mol.

OpenStudy (anonymous):

Wow, thank you so much!(:

OpenStudy (anonymous):

=)